# A unitary vertex operator algebra arising from the $3C$-algebra

### v1

@@ Abstract
We prove that the vertex operator algebra $L(21/22,0)\oplus L(21/22,8)$ is unitary and all its irreducible modules are unitary modules. Moreover, using results from modular tensor categories, we establish a general result about fusion rules for commutant subalgebras under suitable assumptions. As an application, we explicitly determine the fusion rules of all irreducible $L(21/22,0)\oplus L(21/22,8)$-modules.
@@ 1 Introduction
The definition of a unitary vertex operator algebra was first introduced in [6]. Roughly speaking, a vertex operator algebra (VOA for short) is called unitary if it is equipped with a positive definite Hermitian form that is compatible with the vertex operator structure. There are two main motivations for introducing this notion. One arises from its close relationship with unitary representations of infinite dimensional Lie algebras, while the other comes from its importance in connecting the algebraic and analytic approaches to 2-dimensional conformal field theory. We refer the reader to [6] for a detailed discussion. For other approaches to unitary VOAs, see B. Gui’s works [17, 18, 19] and the references therein.
It was shown in [6] that many well-known rational VOAs, such as the unitary series of Virasoro VOAs and lattice VOAs associated to positive definite even lattices, are unitary VOAs. In [7], preunitary VOAs with central charge $c<1$ were classified. Moreover, it was proved there that all such preunitary vertex operator algebras are unitary except for two cases $L(c_{9},0)\oplus L(c_{9},h_{1,7}^{9})$ and $L(c_{27},0)\oplus L(c_{27},h_{1,11}^{27})\oplus L(c_{27},h_{1,19}^{27})\oplus L(c_{27},h_{1,29}^{27}).$ These two VOAs are constructed using the method of mirror extension (see [4, 25]), but there is no uniform approach to establishing the unitarity of VOAs obtained via mirror extension. One of the main goals of the present paper is to establish the unitarity of $L(c_{9},0)\oplus L(c_{9},h_{1,7}^{9}),$ which is isomorphic to $L(21/22,0)\oplus L(21/22,8)$. We now briefly explain the strategy used to achieve this goal. Unlike the realization of $L(21/22,0)\oplus L(21/22,8)$ in [7], which relies mainly on mirror extension theory, our approach is based on coset theory. More precisely, the $3C$-algebra $U_{3C}$ introduced in [24] can be realized as a commutant subalgebra of the lattice vertex operator algebra $V_{\sqrt{2}E_{8}}$. Since lattice vertex operator algebras associated with positive definite even lattices are unitary, it follows from a result in [6] that $U_{3C}$ is unitary. Furthermore, $L(21/22,0)\oplus L(21/22,8)$ can be realized as a commutant subalgebra of $U_{3C}$ and its unitarity can be easily checked. We further show that all irreducible $L(21/22,0)\oplus L(21/22,8)$-modules are unitary.
Another main result of this paper concerns fusion rules of irreducible $L(21/22,0)\oplus L(21/22,8)$-modules. Motivated by the recent work [10], we derive a more general result on fusion rules for commutant subalgebras. More precisely, assume $V$ is a VOA, $U\subset V$ is a subVOA, and $U^{c}$ is the commutant subalgebra of $U$ in $V$, satisfying certain good assumptions (see Section 4 for details). Moreover, we assume all irreducible $U$-modules appear in $V$ when decomposing $V$ into irreducible $U\otimes U^{c}$-modules. Then the fusion rules of $U^{c}$-modules will be determined completely by fusion rules of $V$-modules as well as $U$-modules. As a corollary, we explicitly determine the fusion rules of irreducible $L(21/22,0)\oplus L(21/22,8)$-modules.
@@ Acknowledgments
Xiangyu Jiao is supported by the Science and Technology Commission of Shanghai Municipality (Grant No. 23ZR1418600, No. 22DZ2229014). Wen Zheng is supported by the National Natural Science Foundation of China (No. 12201334) and the Natural Science Foundation of Shandong Province (No. ZR2022QA023).

### v2

@@ Abstract
We give an algebraic proof of the unitarity of the vertex operator algebra $L(21/22,0)\oplus L(21/22,8)$ and of all its irreducible ordinary modules, using a coset realization arising from the $3C$-algebra. Motivated by the structure of the resulting module decomposition, we establish a general result on fusion rules for commutant vertex operator subalgebras within the framework of modular tensor categories. As an application of this general result, we explicitly determine the fusion rules of all irreducible $L(21/22,0)\oplus L(21/22,8)$-modules.
@@ 1 Introduction
The definition of a unitary vertex operator algebra was first introduced in [7]. Roughly speaking, a vertex operator algebra (VOA for short) is called unitary if it is equipped with a positive definite Hermitian form that is compatible with the vertex operator structure. There are two main motivations for introducing this notion. One arises from its close relationship with unitary representations of infinite dimensional Lie algebras, while the second arises from its fundamental role in relating the algebraic formulation of two-dimensional conformal field theory via vertex operator algebras to the analytic approach based on conformal nets. We refer the reader to [7] for a detailed discussion. For other approaches to unitary VOAs, see B. Gui’s works [18, 19, 20, 21] and the references therein.
It was shown in [7] that many well-known rational VOAs, such as the unitary series of Virasoro VOAs and lattice VOAs associated to positive definite even lattices, are unitary VOAs. In [8], preunitary VOAs with central charge $c<1$ were classified, yielding a one to one correspondence with the classification of conformal nets with $c<1$ [23]. In the same work, it was proved there that all such preunitary vertex operator algebras are unitary except for two cases $L(c_{9},0)\oplus L(c_{9},h_{1,7}^{9})$ and $L(c_{27},0)\oplus L(c_{27},h_{1,11}^{27})\oplus L(c_{27},h_{1,19}^{27})\oplus L(c_{27},h_{1,29}^{27})$, whose unitarity remained open from a purely vertex operator algebraic perspective. To bridge these algebraic and analytic classifications, Gui [20] established a fundamental relation between unitary VOA extensions and Q-systems [29] in $C^{*}$-tensor categories. This framework allowed him to prove that the two exceptional VOAs above are indeed unitary, thereby completing the classification of unitary VOAs with $c<1$. See also [2]. Moreover, Gui’s work also shows that all such VOAs are completely unitary [20], which in particular implies that all their irreducible ordinary modules are unitary modules.
The main purpose of the present paper is twofold. First, we give an independent and purely algebraic proof of the unitarity of the vertex operator algebra $L(21/22,0)\oplus L(21/22,8)$ and of all its irreducible ordinary modules. Our approach is based on a concrete coset realization arising from the $3C$-algebra [27]. Second, motivated by a structural feature observed in the explicit decomposition of the $3C$-algebra, we establish a general result on fusion rules for coset vertex operator subalgebras within the framework of modular tensor categories, building in particular on the methods developed in [11].
@@ Acknowledgments
We thank Tiziano Gaudio for his valuable communication and helpful comments.